OpenStudy (anonymous):

Show that sigma of 1/n^2 IS CONVERGES by integral test

hartnn (hartnn):

what are the limits with sigma ?

hartnn (hartnn):

1 to infinity or 2 to infinity or other ?

hartnn (hartnn):

using those limits, try to find \(\int \dfrac{1}{x^2}dx\) if the integral is finite, the sum converges, if not the sum diverges

OpenStudy (anonymous):

to infinity

hartnn (hartnn):

from where to infinity ? 0 or 1 or 2 ?

OpenStudy (anonymous):

from 1

hartnn (hartnn):

ok, then try to solve \(\huge \int \limits_{1}^{\infty} \dfrac{1}{x^2}dx\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

can you solve ? do you know whats integral of x^-2 is ?

OpenStudy (anonymous):

is it 2

hartnn (hartnn):

nopes, how ?

OpenStudy (anonymous):

I confused about that not sure

hartnn (hartnn):

\(\int x^n dx=x^{n+1}/(n+1)+c\) so what about integral of 1/x^2 = x^(-2) ??

OpenStudy (anonymous):

i donot understand

hartnn (hartnn):

have you ever solved an integral ?

OpenStudy (anonymous):

no

hartnn (hartnn):

then why are you using integral test ?

OpenStudy (anonymous):

this new lesson for me

OpenStudy (anonymous):

and i want to understand how it is work

hartnn (hartnn):

i would suggest you to first try out easy examples of integration. then come to this, you will understand only if you have solved some integration questions before.

OpenStudy (anonymous):

i can slove it by other test but i want to understand integral test

hartnn (hartnn):

for that you need to understand integration...

OpenStudy (anonymous):

yes right

hartnn (hartnn):

let me give you one example integral of x^7 will be x^8/8 +c (in x^n, n=7, so n+1 =8) using the formula i gave you above. got this ?

OpenStudy (anonymous):

oh yes i got it

hartnn (hartnn):

so, what will be the integral of x^-2 hint: use n=-2

OpenStudy (anonymous):

x^-1/-1

hartnn (hartnn):

correct! you learn fast :) so, integral of x^-2 is -1/x now put the upper limit in it, that is put x = infinity in -1/x what do u get ?

OpenStudy (anonymous):

do you mean x^-1/-1 i have to put x= infinity

OpenStudy (anonymous):

or just for -1/x

hartnn (hartnn):

yes. the idea is . \(\int \limits_{lower \: limit }^{upper \: limit }f(x)dx = F(upper \:limit)-F(lower \: limit)\) and x^-1/-1 and -1/x are SAME thing.

hartnn (hartnn):

so put the upper limit in it, that is put x = infinity in -1/x what do u get ?

OpenStudy (anonymous):

0 =0

hartnn (hartnn):

good, now put the lower limit in it, that is put x = 1 in -1/x what do u get ?

OpenStudy (anonymous):

=-1

hartnn (hartnn):

correct, so we have \(\huge \int \limits_{1}^{\infty} \dfrac{1}{x^2}dx= -1/x|_{x=1}^{x=\infty}= 0-(-1)=1\) got this ? so we got the integral as FINITE. what does we conclude from this ?

OpenStudy (anonymous):

yes i got it now

hartnn (hartnn):

integral FINITE means the sum converges :)

OpenStudy (anonymous):

if its equal to one or lease than one mean converges

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i have other question

OpenStudy (anonymous):

thank you so much

hartnn (hartnn):

sorry, i was away from keyboard, welcome ^_^ whats the other question ?

hartnn (hartnn):

and for convergence by integral test, we see whether the answer is FINITE or not, we don't compare it with 1 if finite, then converges